Operations Report – January 28th

Matt Wise
MDRS Crew 220

Operations Report

Crew 220 Operations Report 28JAN2020

Sol: 9

Name of person filing report: Matt Wise

Non-nominal systems: Primary Generator, Water Heater, Hab lower deck fluorescent lights

Generator: Primary is inoperable. Rental was run-the following information is for the rental.

Hours run: 12hrs
From what time last night: 21:00

To what time this morning: 08:00

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 81%

Diesel Reading – 1/3

Station Propane Reading- 81

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 6

Water Meter (units)- 0148436,0

Water (static tank) (gal) – 161

Static to loft Pump used – Yes 40gal

Water in GreenHab (gal): 136.64

Water in Science Dome (gal): 0

Toilet Tank Emptied – No

Deimos rover used: Still in the workshop


Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: No


Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used – Yes

Starting Hours: 128

Beginning Charge: 100%

Ending Hours: 128.3

Ending Charge: 62%

Opportunity Rover used: Still in workshop


Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

Starting Hours: 134

Beginning Charge: 100%

Ending Hours: 134.3

Ending Charge: 63%

Notes on rovers: Opportunity and Deimos off-site for maintenance. There were no EVAs today, so the rovers were not used.

ATVs Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today: 0

Notes on ATVs: None.

Habcar used and why, where? No

CrewCar used and why, where? No

General Notes and Comments: None.

Summary of the internet: Nominal

Summary of Suit and Radios: 3 radios malfunctioning
Suits Nominal

Summary of Hab: Main fluorescent lights in Hab lower deck malfunctioning. Center light bar now dim and flickering all times, light bar near back airlock not functioning, and light bar near stairs has no bulbs. Engineer attempted to remove/replace bulbs with no effect. We’re not aware of any replacement bulbs in Hab or RAM and will be requesting bulbs in the next resupply from Outpost.

Summary of Science Dome operations: Nominal

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and Safety issues: MDRS crew all healthy

Questions, concerns, and requests to Mission Support: We’d like to request 4x fluorescent bulbs in the next resupply. Thanks!

Copyright © The Mars Society. All rights reserved. | Main Site